Twice Sally's age increased by 3 times Sue's age is 103. Five times Sally's age decreased by 4 times Sue's age is 16. How old will Sally be ten years from now?
Answer by ankor@dixie-net.com(22740)

Twice Sally's age increased by 3 times Sue's age is 103.
Five times Sally's age decreased by 4 times Sue's age is 16.
How old will Sally be ten years from now?
:
Let x = Sally's present age
Let y = Sue's present age
:
Write an equation for each statement:
"Twice Sally's age increased by 3 times Sue's age is 103."
2x + 3y = 103
:
"Five times Sally's age decreased by 4 times Sue's age is 16."
5x - 4y = 16
:
To use elimination here, multiply the 1st equation by 4, the 2nd by 3:
8x + 12y = 412
15x -12y = 48
------------------addition eliminates y, find x
23x = 460
x = 460/23
x = 20 yr is Sally's present age
:
" How old will Sally be ten years from now?"
I'll let you figure that out